OpenStudy (anonymous):

Jason considered two similar televisions at a local electronics store. The generic version was based off the brand name and was the size of the brand name. If the generic television set is 12 inches by 24 inches, what are the dimensions of the brand name television?

OpenStudy (anonymous):

18 inches by 36 inches 6 inches by 12 inches 16 inches by 32 inches 9 inches by 18 inches

OpenStudy (anonymous):

Help @Fruitbasket !?

OpenStudy (anonymous):

You missed out how much larger/smaller the generic version is.

OpenStudy (anonymous):

The generic version was based off the brand name and was 3/4 the size of the brand name.

OpenStudy (anonymous):

Okay. In that case, the dimensions of the brand name television are the dimensions of the generic television, just divided by 3/4. So the dimensions will be 12/(3/4) = (12 * 4)/3 = 16 inches by 24/(3/4) = (24*4)/3 = 32 inches.

OpenStudy (anonymous):

Thank you !